# AiXpand AI OS - Decentralized ubiquitous computing MLOps execution engine

### v1

@@ 1 Introduction
1.2 The AiXpand vision
@@ 1.3 Summarizing our progress so far
In our previous work [11], we proposed an end-to-end architecture and methodology aimed at standardizing critical stages of production-grade machine learning pipelines, with a particular focus on edge-based systems. The primary goal was to offer technological independence, freedom to operate, and versatility in composing and deploying business rules. This resulted in an end-to-end MLOps framework that supports multiple jobs and workers, tensor framework agnosticism, and pre-built templates and serving pipelines for prominent frameworks such as TensorFlow [13] and PyTorch [14]. The currently proposed architecture enables seamless integration with a wide array of data sources, including CCTV cameras, relational databases, flat files, and sensor networks, while facilitating rapid deployment of business rules through low-code and no-code approaches. The component which lies at the foundation of this architecture, the AiXpand Execution Engine, can be easily configured for IoT-based protocols like MQTT [15] and AMQP [16], as well as web REST endpoints, with other endpoints implementable via a low-code plugin approach.
@@ 2.1 MLOps, DevOps and productization of AI
Finally, in relation to the AiXpand architecture, while the above-mentioned frameworks and technologies could be considered competing approaches for MLOps, most of them are in fact technologies that could be easily integrated and encapsulated within the AiXpand Execution Engine Pipelines.
@@ 2.2 From edge, fog & cloud to AI
Currently, there are three main approaches to providing AI-enabled applications: edge, fog, and cloud AI. Edge AI [21] involves deploying AI applications directly on the devices where data is generated. This approach allows users to utilize moderately advanced hardware solutions that are in close proximity to the data that requires processing. It is a cost-effective method that offers enhanced privacy. However, it lacks scalability when there is a growing demand for AI solutions. On the other hand, cloud AI enables vertical scaling by leveraging increasing computing power as needed. Nevertheless, it faces limitations at the network level when dealing with substantial amounts of data. Fog AI offers a horizontal scaling solution by utilizing a centralized network of edge devices. It aims to fulfill the high computing requirements of cloud AI while reducing overall network traffic. Both fog and cloud AI adopt subscription-based business models, which may not be advantageous for businesses that do not require real-time AI solutions. We propose a decentralized architecture that harnesses the computing capabilities of edge devices while maintaining low-latency connections between customers and AiXpand network nodes. Our architecture provides an on-demand, micro-transaction payment-based option for clients with AI-powered end-to-end application needs.
@@ 3.1 A top-down view
The proposed decentralized architecture, as illustrated in Figure1, is grounded in the well-established design of validator-like AiXpand Processing Nodes, which facilitate regional or global aggregation of AiXpand Processing Node Pools. Distinct from traditional validators responsible for blockchain protocol consensus, the primary function of an AiXpand Validator Node pertains to job validation and communication aggregation. Concurrently, all AiXpand Processing Nodes participate in a private blockchain consensus, ensuring secure and immutable job execution and data transfers.
Figure 1: Overall view of the AiXpand E2s network of networks
The overall AiXpand Ecosystem encompasses various middleware components, including IoT message broker tools and REST APIs for diverse use-cases. At its core is the AiXpand Execution Engine (E2), depicted in Figure 2, along with its platform-agnostic SDKs. The E2 can be deployed on a wide array of target operating systems and hardware devices, with or without advanced GPU compute capabilities. AiXpand employs both a private proprietary lite blockchain for internal use and an external public mainnet blockchain for microtransactions orchestrated by individual E2 processing nodes. Consequently, each E2 is uniquely identified and recognized within the network through a public blockchain (mainnet) non-fungible utility token. The subsequent paragraphs will discuss the primary decentralized processing distribution and the integration of a microtransaction-based worker reward system.
Although a comprehensive tokenomics strategy is proposed, it has not yet been publicly released. This paper focuses on the high-level technological aspects related to the AiXpand E2 Node Deeds’ interaction with the protocol, the fundamental principles of oracle-like communication between the mainnet blockchain (BC) and AiXpand E2s, and the overall algorithmic approach for reward distribution 321.
Figure 2: Top view of AiXpand Execution Engine ecosystem
@@
The Proof-of-AI protocol (PoAI) 1 manages the rewarding of users who operate AiXpand E2 processing nodes. This incentivizes the provision of compute power by both miners and service consumers. Miners use the AiXpand E2 solely for decentralized job execution, while service consumers use specific functionalities or applications developed by AiXpand content creators and service providers. As previously mentioned, an in-depth explanation of public on-chain algorithms and procedures is beyond the purpose of this paper, however the pseudo-code algorithmic approach of AiXpand Proof of AI (PoAI) can be analyzed in the current paper.
In order to summarize the funds allocation and distribution in our job execution agnostic reward pools we will denote: $T_{N}$ the total time a given node $N$, identified by a unique smart-contract based on a Node Deed in the overall population of all $A$ AiXpand Processing Nodes, is alive in the network; $T_{epoch}$ the protocol epoch time which is the interval used to calculate and distribute a round of rewards to all active nodes; $E$ the total number of epochs since the protocol genesys; $P_{N_{E}}$ the power score of a given node $N$ during epoch $E$. The final formula of the reward $R_{N_{E}}$ of the $Node$ $N$ at epoch $E$ is presented below in Equation5.
@@ 3.2 End-to-end low-code pipelines
The AiXpand Ecosystem aims to simplify the process of developing and deploying complex application systems by dividing it into four distinct heterogeneous subsystems: communication, data acquisition, model serving, and business process as seen in Figure2. Each subsystem includes its own highly composable plugins, each with configurable parameters that allow customization of functionality without requiring additional code to be written. This versatility is achieved by allowing a pipeline to be configured by specifying which plugins to use and how they should be configured, enabling easy alteration of the pipeline’s functionality without code intervention or with an extremely low code footprint. This structure allows developers to use pre-existing or create custom low-code solutions for each of the subsystem within the network’s ecosystem, simplifying development, deployment and integration of complex application systems.
@@ 3.3 Secure execution with Plugin API
Two primary types of business plugins have been developed, namely secure plugins and user plugins. These plugins are currently built using Python[33], a high-level general-purpose programming language. Business plugins can be considered either generic templates or default plugins for preparing business insights data/reports. On the other hand, secure plugins encompass a range of user plugins designed as proposals for future secure services, as well as code snippets created by anyone using the AiXpand E2 within the trustless network.
@@
It is important to note that all these approaches can be accessed and utilized through external endpoints. This can be done either via a JavaScript/TypeScript-based SDK or through a Python client that utilizes our AiXpand E2 PyE2 package.
@@ 3.4 Multi-framework serving processes
The AiXpand Network and its ecosystem are designed to support applications powered AI. This is achieved by incorporating various AI models through the use of AI serving plugins within the serving subsystem. Additionally, the ecosystem is designed to accommodate applications that do not rely on AI models or techniques but are well-suited for decentralized environments.
To enable this, AiXpand provides a serving plugin interface, illustrated in Figure 3, which is not limited to specific frameworks like TensorFlow [13]. In this setup, all data input, inference, and prediction output are handled within the main thread of the main process. The implementation of the serving plugin interface primarily involves model loading, data pre-processing, and inference post-processing.
@@
Figure 3: High-level AiXpand E2 model serving processes UML diagram
@@ 3.5 Distributed decentralized execution
The architecture of the AiXpand Network is specifically designed to achieve scalable and decentralized distribution of different processing workloads. This is accomplished through a combination of peer-to-peer trustless job execution based on Map Reduce [17] and Scatter Gather [34] techniques, along with DLT. Throughout the design process, ensuring the security of all participants has been of paramount importance, influencing the network architecture and the distribution of processing workloads.
@@
To provide a concise overview of the high-level process, we define an instance of the AiXpand E2 that requires external processing offload as a Main Node, while other nearby AiXpand E2 instances will be referred to as potential Worker Nodes. In this context, the Main Nodes have the capability to initiate distributed execution pipelines, each comprising two main processing stages, each with two distinct steps.
@@
The primary goal of the AiXpand Ecosystem is to cater to a broad range of participants, including content creators, processing node operators, software developers, and end-consumers. To achieve this objective, AiXpand provides a comprehensive platform that is adaptable, reliable, secure, and easily scalable. This platform is designed to support not only AI-powered applications but also other decentralized applications, regardless of whether they utilize AI models or not.
@@
Through its compatibility with various frameworks and the ability to execute multiple model-serving processes in parallel, AiXpand ensures an efficient and flexible solution for its users. The platform’s distinctive design, coupled with its integration with blockchain technology, enables secure and transparent operations, effectively meeting the diverse needs of participants.
@@
To enhance the adoption of the ecosystem by the public, software development kits (SDKs) have been developed. These SDKs support multiple programming languages and frameworks, such as Python and NodeJS, significantly expanding integration options and improving the user experience. By using the SDKs, users gain access to the AiXpand Network through a partially functional paradigm, which offers callbacks for manipulating payloads received from AiXpand Nodes and provides simple yet powerful features for communication with the network. This empowers users to connect to the network, establish new end-to-end pipelines or join existing ones, create predefined or custom jobs within the pipeline, and receive notifications and heartbeats, facilitating seamless interaction with the ecosystem.
@@
The AiXpand SDKs are intended to be used as a cohesive integration tool between the AI processing network and any existing or new software applications. They are designed to automate and manage the lifecycle of requests made on the AiXpand Network. AiXpand SDKs come in several flavors targeting either the Python or Node.js developers’ communities.
@@
In the annex of this paper, we present a simple Python 3.x example 1 that demonstrates the process of connecting to a processing AiXpand E2 Node, executing tasks remotely, and distributing an end-to-end pipeline. However, it’s important to note that this example does not include the necessary steps to enable Proof-of-AI (PoAI) reward distribution, which involves compensating the processing node pool for the executed job.
3.5.1 The AiXpand SDKs
@@
The AiXpand Node.js SDK consists of two main components. The first component aids in project setup and maintenance, handling tasks such as code generation and message validation using YAML schemas. The second component is responsible for runtime operations, serving as an intermediary for communication with the network and an example of it can be seen in example 2
@@
To support developers in integrating network functionality into their solutions, a Command Line Interface (CLI) tool is provided. This tool assists in constructing the necessary schemas for communicating with deployed pipelines in the AiXpand Network. During project development or maintenance, developers can use the CLI to ensure that the command definitions for the pipelines are synchronized with the network.
@@
To incorporate AiXpand into a project, there are two approaches available. The first option is to execute a series of CLI commands using the AiXpand Node.js SDK. This will retrieve the necessary AI pipeline message schemas specifically designed for Node.js applications. Alternatively, for Python development teams, the AiXpand PyE2 SDK package can be installed, eliminating the need for CLI commands. Once this prerequisite is fulfilled, the imported libraries within the project will handle all validation and transformations between the application’s business objects and the expected message formats of the AiXpand Nodes.
To publish a message on the network, it is essential to establish a valid communication channel with a public AiXpand MQTT. This can be achieved by connecting to the AiXpand Validator Node or by using the local broker bundled with the AiXpand Node. Both options provide extensive availability of resources and published AI services. Furthermore, AiXpand PyE2 SDK provides built-in support for listening to payloads, intercepting errors, and diagnosing any issues that may arise due to incorrect use of the AI pipeline plugins.
@@ 3.6 Integration with Web3
The use of on-demand AI functionalities and capabilities offered by the AiXpand Network involves two main stages. The first stage involves the financial commitment of the processing power consumer, where a designated amount of tokens is allocated for covering AI request-related expenses. In the second stage, decentralized job execution takes place.
@@
To facilitate efficient communication between local nodes, the AiXpand Network employs the AiXpand communication protocol, which utilizes MQTT as the underlying transport layer. MQTT is a widely adopted protocol known for its reliable scalability in IoT applications. The lifecycle of the AiXpand Protocol is encapsulated within the AiXpand PyE2 SDK, simplifying its implementation and usage.
@@
To guarantee compensation for potential worker nodes’ efforts, it is essential that network requests are accompanied by a financial commitment. By including information pertaining to a valid inbound AiXpand wallet transaction, an AiXpand Node can process a designated workload. This step can be executed by the consumer application or via a separate procedure, as long as it is completed before submitting the workload to the network.
@@
The AiXpand PyE2 SDK’s primary roles are to manage relevant financial information before disseminating the request on the network and to simplify the integration of this technology within external systems. For processes that necessitate multiple network requests, substantial transaction fees may be incurred. To minimize expenses related to utilizing cryptocurrency networks, depositing a larger sum into an escrow account is possible, allowing the AiXpand internal blockchain to oversee expended funds.
@@
The AiXpand PyE2 SDK employs specific Web3 components for interacting with payment channels, contingent on implementation. On the AiXpand E2 side, a simple validation is performed to determine whether the request is signed by the AiXpand Node owner and run locally (in which case payment is unnecessary) or whether a valid funding transaction is attached.
@@
The request entails disseminating a simple message to a user-controlled node or the nearest validator node. Subsequently, the network assumes responsibility for managing the request. Built with robust components, the AiXpand PyE2 SDK offers a reliable communication channel with the AI-enabled node network. Request progression is monitored by event listeners within the SDK, and any outcome is documented and relayed to the initiating processes.
@@
Throughout pipeline execution, AiXpand Network Nodes generate status payloads, which outline the request’s current state, report potential errors, and supply information on generated assets. These payloads are monitored by event listeners within the AiXpand PyE2 SDK components, and appropriate handlers are provided for seamless interaction with ongoing processes.
3.7 AiXpand Rapid Application Development

### v2

# Naeural AI OS - Decentralized ubiquitous computing MLOps execution engine

@@ 1 Introduction
1.2 The Naeural Edge Protocol vision
@@ 1.3 Summarizing our progress so far
In our previous work [11], we proposed an end-to-end architecture and methodology aimed at standardizing critical stages of production-grade machine learning pipelines, with a particular focus on edge-based systems. The primary goal was to offer technological independence, freedom to operate, and versatility in composing and deploying business rules. This resulted in an end-to-end MLOps framework that supports multiple jobs and workers, tensor framework agnosticism, and pre-built templates and serving pipelines for prominent frameworks such as TensorFlow [13] and PyTorch [14]. The currently proposed architecture enables seamless integration with a wide array of data sources, including CCTV cameras, relational databases, flat files, and sensor networks, while facilitating rapid deployment of business rules through low-code and no-code approaches. The component which lies at the foundation of this architecture, the Naeural Execution Engine, can be easily configured for IoT-based protocols like MQTT [15] and AMQP [16], as well as web REST endpoints, with other endpoints implementable via a low-code plugin approach.
@@ 2.1 MLOps, DevOps and productization of AI
Finally, in relation to the Naeural architecture, while the above-mentioned frameworks and technologies could be considered competing approaches for MLOps, most of them are in fact technologies that could be easily integrated and encapsulated within the Naeural Execution Engine Pipelines.
@@ 2.2 From edge, fog & cloud to AI
Currently, there are three main approaches to providing AI-enabled applications: edge, fog, and cloud AI. Edge AI [21] involves deploying AI applications directly on the devices where data is generated. This approach allows users to utilize moderately advanced hardware solutions that are in close proximity to the data that requires processing. It is a cost-effective method that offers enhanced privacy. However, it lacks scalability when there is a growing demand for AI solutions. On the other hand, cloud AI enables vertical scaling by leveraging increasing computing power as needed. Nevertheless, it faces limitations at the network level when dealing with substantial amounts of data. Fog AI offers a horizontal scaling solution by utilizing a centralized network of edge devices. It aims to fulfill the high computing requirements of cloud AI while reducing overall network traffic. Both fog and cloud AI adopt subscription-based business models, which may not be advantageous for businesses that do not require real-time AI solutions. We propose a decentralized architecture that harnesses the computing capabilities of edge devices while maintaining low-latency connections between customers and Naeural network nodes. Our architecture provides an on-demand, micro-transaction payment-based option for clients with AI-powered end-to-end application needs.
@@ 3.1 A top-down view
The proposed decentralized architecture, as illustrated in Figure1, is grounded in the well-established design of validator-like Naeural Processing Nodes, which facilitate regional or global aggregation of Naeural Processing Node Pools. Distinct from traditional validators responsible for blockchain protocol consensus, the primary function of an Naeural Validator Node pertains to job validation and communication aggregation. Concurrently, all Naeural Processing Nodes participate in a private blockchain consensus, ensuring secure and immutable job execution and data transfers.
Figure 1: Overall view of the Naeural E2s network of networks
The overall Naeural Ecosystem encompasses various middleware components, including IoT message broker tools and REST APIs for diverse use-cases. At its core is the Naeural Execution Engine (E2), depicted in Figure 2, along with its platform-agnostic SDKs. The E2 can be deployed on a wide array of target operating systems and hardware devices, with or without advanced GPU compute capabilities. Naeural employs both a private proprietary lite blockchain for internal use and an external public mainnet blockchain for microtransactions orchestrated by individual E2 processing nodes. Consequently, each E2 is uniquely identified and recognized within the network through a public blockchain (mainnet) non-fungible utility token. The subsequent paragraphs will discuss the primary decentralized processing distribution and the integration of a microtransaction-based worker reward system.
Although a comprehensive tokenomics strategy is proposed, it has not yet been publicly released. This paper focuses on the high-level technological aspects related to the Naeural E2 Node Deeds’ interaction with the protocol, the fundamental principles of oracle-like communication between the mainnet blockchain (BC) and Naeural E2s, and the overall algorithmic approach for reward distribution 321.
Figure 2: Top view of Naeural Execution Engine ecosystem
@@
The Proof-of-AI protocol (PoAI) 1 manages the rewarding of users who operate Naeural E2 processing nodes. This incentivizes the provision of compute power by both miners and service consumers. Miners use the Naeural E2 solely for decentralized job execution, while service consumers use specific functionalities or applications developed by Naeural content creators and service providers. As previously mentioned, an in-depth explanation of public on-chain algorithms and procedures is beyond the purpose of this paper, however the pseudo-code algorithmic approach of Naeural Proof of AI (PoAI) can be analyzed in the current paper.
In order to summarize the funds allocation and distribution in our job execution agnostic reward pools we will denote: $T_{N}$ the total time a given node $N$, identified by a unique smart-contract based on a Node Deed in the overall population of all $A$ Naeural Processing Nodes, is alive in the network; $T_{epoch}$ the protocol epoch time which is the interval used to calculate and distribute a round of rewards to all active nodes; $E$ the total number of epochs since the protocol genesys; $P_{N_{E}}$ the power score of a given node $N$ during epoch $E$. The final formula of the reward $R_{N_{E}}$ of the $Node$ $N$ at epoch $E$ is presented below in Equation5.
@@ 3.2 End-to-end low-code pipelines
The Naeural Ecosystem aims to simplify the process of developing and deploying complex application systems by dividing it into four distinct heterogeneous subsystems: communication, data acquisition, model serving, and business process as seen in Figure2. Each subsystem includes its own highly composable plugins, each with configurable parameters that allow customization of functionality without requiring additional code to be written. This versatility is achieved by allowing a pipeline to be configured by specifying which plugins to use and how they should be configured, enabling easy alteration of the pipeline’s functionality without code intervention or with an extremely low code footprint. This structure allows developers to use pre-existing or create custom low-code solutions for each of the subsystem within the network’s ecosystem, simplifying development, deployment and integration of complex application systems.
@@ 3.3 Secure execution with Plugin API
Two primary types of business plugins have been developed, namely secure plugins and user plugins. These plugins are currently built using Python[33], a high-level general-purpose programming language. Business plugins can be considered either generic templates or default plugins for preparing business insights data/reports. On the other hand, secure plugins encompass a range of user plugins designed as proposals for future secure services, as well as code snippets created by anyone using the Naeural E2 within the trustless network.
@@
It is important to note that all these approaches can be accessed and utilized through external endpoints. This can be done either via a JavaScript/TypeScript-based SDK or through a Python client that utilizes our Naeural E2 PyE2 package.
@@ 3.4 Multi-framework serving processes
The Naeural Network and its ecosystem are designed to support applications powered AI. This is achieved by incorporating various AI models through the use of AI serving plugins within the serving subsystem. Additionally, the ecosystem is designed to accommodate applications that do not rely on AI models or techniques but are well-suited for decentralized environments.
To enable this, Naeural provides a serving plugin interface, illustrated in Figure 3, which is not limited to specific frameworks like TensorFlow [13]. In this setup, all data input, inference, and prediction output are handled within the main thread of the main process. The implementation of the serving plugin interface primarily involves model loading, data pre-processing, and inference post-processing.
@@
Figure 3: High-level Naeural E2 model serving processes UML diagram
@@ 3.5 Distributed decentralized execution
The architecture of the Naeural Network is specifically designed to achieve scalable and decentralized distribution of different processing workloads. This is accomplished through a combination of peer-to-peer trustless job execution based on Map Reduce [17] and Scatter Gather [34] techniques, along with DLT. Throughout the design process, ensuring the security of all participants has been of paramount importance, influencing the network architecture and the distribution of processing workloads.
@@
To provide a concise overview of the high-level process, we define an instance of the Naeural E2 that requires external processing offload as a Main Node, while other nearby Naeural E2 instances will be referred to as potential Worker Nodes. In this context, the Main Nodes have the capability to initiate distributed execution pipelines, each comprising two main processing stages, each with two distinct steps.
@@
The primary goal of the Naeural Ecosystem is to cater to a broad range of participants, including content creators, processing node operators, software developers, and end-consumers. To achieve this objective, Naeural provides a comprehensive platform that is adaptable, reliable, secure, and easily scalable. This platform is designed to support not only AI-powered applications but also other decentralized applications, regardless of whether they utilize AI models or not.
@@
Through its compatibility with various frameworks and the ability to execute multiple model-serving processes in parallel, Naeural ensures an efficient and flexible solution for its users. The platform’s distinctive design, coupled with its integration with blockchain technology, enables secure and transparent operations, effectively meeting the diverse needs of participants.
@@
To enhance the adoption of the ecosystem by the public, software development kits (SDKs) have been developed. These SDKs support multiple programming languages and frameworks, such as Python and NodeJS, significantly expanding integration options and improving the user experience. By using the SDKs, users gain access to the Naeural Network through a partially functional paradigm, which offers callbacks for manipulating payloads received from Naeural Nodes and provides simple yet powerful features for communication with the network. This empowers users to connect to the network, establish new end-to-end pipelines or join existing ones, create predefined or custom jobs within the pipeline, and receive notifications and heartbeats, facilitating seamless interaction with the ecosystem.
@@
The Naeural SDKs are intended to be used as a cohesive integration tool between the AI processing network and any existing or new software applications. They are designed to automate and manage the lifecycle of requests made on the Naeural Network. Naeural SDKs come in several flavors targeting either the Python or Node.js developers’ communities.
@@
In the annex of this paper, we present a simple Python 3.x example 1 that demonstrates the process of connecting to a processing Naeural E2 Node, executing tasks remotely, and distributing an end-to-end pipeline. However, it’s important to note that this example does not include the necessary steps to enable Proof-of-AI (PoAI) reward distribution, which involves compensating the processing node pool for the executed job.
3.5.1 The Naeural SDKs
@@
The Naeural Node.js SDK consists of two main components. The first component aids in project setup and maintenance, handling tasks such as code generation and message validation using YAML schemas. The second component is responsible for runtime operations, serving as an intermediary for communication with the network and an example of it can be seen in example 2
@@
To support developers in integrating network functionality into their solutions, a Command Line Interface (CLI) tool is provided. This tool assists in constructing the necessary schemas for communicating with deployed pipelines in the Naeural Network. During project development or maintenance, developers can use the CLI to ensure that the command definitions for the pipelines are synchronized with the network.
@@
To incorporate Naeural into a project, there are two approaches available. The first option is to execute a series of CLI commands using the Naeural Node.js SDK. This will retrieve the necessary AI pipeline message schemas specifically designed for Node.js applications. Alternatively, for Python development teams, the Naeural PyE2 SDK package can be installed, eliminating the need for CLI commands. Once this prerequisite is fulfilled, the imported libraries within the project will handle all validation and transformations between the application’s business objects and the expected message formats of the Naeural Nodes.
To publish a message on the network, it is essential to establish a valid communication channel with a public Naeural MQTT. This can be achieved by connecting to the Naeural Validator Node or by using the local broker bundled with the Naeural Node. Both options provide extensive availability of resources and published AI services. Furthermore, Naeural PyE2 SDK provides built-in support for listening to payloads, intercepting errors, and diagnosing any issues that may arise due to incorrect use of the AI pipeline plugins.
@@ 3.6 Integration with Web3
The use of on-demand AI functionalities and capabilities offered by the Naeural Network involves two main stages. The first stage involves the financial commitment of the processing power consumer, where a designated amount of tokens is allocated for covering AI request-related expenses. In the second stage, decentralized job execution takes place.
@@
To facilitate efficient communication between local nodes, the Naeural Network employs the Naeural communication protocol, which utilizes MQTT as the underlying transport layer. MQTT is a widely adopted protocol known for its reliable scalability in IoT applications. The lifecycle of the Naeural Protocol is encapsulated within the Naeural PyE2 SDK, simplifying its implementation and usage.
@@
To guarantee compensation for potential worker nodes’ efforts, it is essential that network requests are accompanied by a financial commitment. By including information pertaining to a valid inbound Naeural wallet transaction, an Naeural Node can process a designated workload. This step can be executed by the consumer application or via a separate procedure, as long as it is completed before submitting the workload to the network.
@@
The Naeural PyE2 SDK’s primary roles are to manage relevant financial information before disseminating the request on the network and to simplify the integration of this technology within external systems. For processes that necessitate multiple network requests, substantial transaction fees may be incurred. To minimize expenses related to utilizing cryptocurrency networks, depositing a larger sum into an escrow account is possible, allowing the Naeural internal blockchain to oversee expended funds.
@@
The Naeural PyE2 SDK employs specific Web3 components for interacting with payment channels, contingent on implementation. On the Naeural E2 side, a simple validation is performed to determine whether the request is signed by the Naeural Node owner and run locally (in which case payment is unnecessary) or whether a valid funding transaction is attached.
@@
The request entails disseminating a simple message to a user-controlled node or the nearest validator node. Subsequently, the network assumes responsibility for managing the request. Built with robust components, the Naeural PyE2 SDK offers a reliable communication channel with the AI-enabled node network. Request progression is monitored by event listeners within the SDK, and any outcome is documented and relayed to the initiating processes.
@@
Throughout pipeline execution, Naeural Network Nodes generate status payloads, which outline the request’s current state, report potential errors, and supply information on generated assets. These payloads are monitored by event listeners within the Naeural PyE2 SDK components, and appropriate handlers are provided for seamless interaction with ongoing processes.
3.7 Naeural Rapid Application Development